# Verifiable Fine-Tuning for LLMs: Zero-Knowledge Training Proofs Bound to Data Provenance and Policy

### v1

@@ I Introduction
Large language models (LLMs) have rapidly migrated from research prototypes to critical infrastructure for search, assistance, and decision support, yet the trust gap around how these models are trained and fine-tuned is widening. Concerns span from whether proprietary or sensitive data were used, to whether claimed training configurations actually produced the released checkpoints, and to whether post-hoc forensics can establish accountability when things go wrong [2]. Conventional privacy and security analyses underscore that inferring whether specific records participated in training is fundamentally limited, which makes ex post verification particularly tenuous [23, 24]. Proof-of-learning (PoL) pushes the opposite direction by seeking cryptographic or audit-style evidence that the training trajectory is genuine [1], and recent variants address incentive compatibility when proofs drive economic or governance consequences [19].
@@
A parallel thread pursues verifiable machine learning (VML), where the computation itself is accompanied by succinct proofs. Early systems established feasibility for neural inference under zero-knowledge (ZK) or related proof systems [3, 5, 15]. More recently, researchers have begun to lift verification from inference to (parts of) training by circuitizing differentiable updates and optimizer states [9, 11]. For LLMs, architectural regularities and approximation-friendly algebra enable practical proof composition, suggesting a path toward end-to-end guarantees in realistic instruction-tuning pipelines [13]. Methodologically, advances in proof-friendly tensor approximations (e.g., polynomial constraints over softmax) and series-expansion techniques further reduce prover overheads and widen coverage [14]. Survey and position papers consolidate these advances and highlight open challenges around scalability, recursion, and on-chain verification for ML [12, 6, 8].
@@
Beyond cryptographic protocols, the surrounding audit stack must establish that the data and configurations claimed for training are accurate. Dataset-level attestation proposes to certify distributional properties without revealing examples [7], while data-engineering provenance offers pragmatic hooks by capturing end-to-end lineage across preparation and experimentation workflows [4, 10]. Systems work is beginning to close the loop by treating provenance as a first-class runtime artifact amenable to programmatic enforcement and reproducibility [35], and by exploring how LLMs themselves can assist in provenance inference when metadata are incomplete [25]. Complementarily, model-provenance testing proposes behavioral probes to ascribe models to sources or families when direct logs are unavailable [16]. Together with PoL-style arguments [1, 19], these strands motivate holistic, multi-layer verification.
@@
Watermarking techniques have also been advanced as a lightweight provenance signal for model outputs, but their reliability is contested. Classic and modern schemes embed statistical beacons into generated text and images [26, 33], with recent work demonstrating scalable detectors for LLM text at industrial scale [27]. Yet watermarks can be fragile or learnable by downstream models [28], susceptible to removal or theft [29, 32], and often deliver weakest guarantees precisely where high-stakes provenance is needed, prompting more robust coding-theoretic designs [30, 31] and broader surveys that map the design space and limitations [34]. These observations reinforce the case for training-time verifiability that does not rely solely on output beacons.
@@
In parallel, trusted execution and attestation are being used to produce verifiable “property cards” about ML artifacts, offering complementary hardware-backed assurances when cryptographic proofs are too costly or incomplete [21]. However, TEEs alone cannot resolve disputes about which data were used or whether the optimizer trajectory conformed to policy, again pointing toward hybrid designs that combine attestations, provenance capture, and cryptographic proofs.
@@
The need for verifiable training is especially acute in federated and edge-centric learning, where heterogeneous clients, intermittent connectivity, and privacy constraints complicate governance. Practical systems already struggle with non-IID data, communication bottlenecks, and asynchrony [37, 59, 41]. Algorithmic advances target adaptive regularization and knowledge distillation for robustness [45], decentralized aggregation with layer-wise strategies [50], and personalization frameworks that better reflect client idiosyncrasies [43, 47]. Neural architecture search and progressive training mitigate heterogeneity in on-device settings [44, 52], while LoRA- and quantization-aware fine-tuning reduce resource footprints for practical deployments [54]. Recent work explores probabilistic communication and local-update strategies to accelerate decentralized training under diverse edges [60, 61], pipelines for collaborative end–cloud inference [64], strategies for MoE inference efficiency [62], communication-efficient decentralized graph learning [58], and mitigating catastrophic forgetting during federated fine-tuning [63]. These directions improve scalability but simultaneously enlarge the attack surface and the number of untrusted actors—prime territory for verifiable fine-tuning protocols that certify data usage and update correctness [35, 6].
@@
Ubiquitous sensing further complicates the risk calculus. Wi-Fi and RF modalities, often intertwined with intelligent assistants, have been shown to support keystroke eavesdropping from commodity microphones [38], undetectable attacks on PHY-layer fingerprint-based authentication [36], and open-set gesture and attention-based recognition that blur lines between ambient intelligence and surveillance [40, 56]. Multimodal affective computing demonstrates how Wi-Fi CSI and vision together can unobtrusively infer emotion, expanding the privacy stakes of data pipelines [48]. On the defensive and robust-sensing side, prior work models and mitigates co-channel interference in activity recognition through subcarrier correlation selection and interference-independent phase components [51, 49]. Healthcare-oriented Wi-Fi sensing shows both potential and sensitivity, from target-oriented respiratory monitoring to pulmonary function estimation without physical contact [53, 55]. RF-based writing and identity inference illustrates how ordinary RFID can leak highly personal behaviors [57]. Even adjacent work on label-noise suppression and distributionally robust facial expression recognition underscores how data curation and reliability shape downstream behaviors that later demand auditing [42, 65]. These examples motivate verifiable training pipelines that can convincingly demonstrate which signals were included, which were excluded, and how updates adhered to declared policies [4].
@@
Our perspective in this paper is that verifiable fine-tuning should be treated as a first-class systems problem that composes three pillars: (i) cryptographic protocols that prove the existence of a legitimate gradient-based path from a public initialization to the released model under declared hyperparameters [9, 11, 13], (ii) provenance capture that binds those proofs to accountable, inspectable data and configuration claims [10, 35, 16], and (iii) audit mechanisms that recognize the limits of post-hoc inference and instead emphasize pre-commitment and incentive-compatible verification [1, 18, 19]. We position our approach within this emerging landscape and focus on end-to-end verifiability for fine-tuning, complementing watermarking-based output attribution [26, 27] and hardware-assisted attestations [21], while remaining deployable in federated and edge scenarios [44, 43].
@@ II-A From Proof-of-Learning to Verifiable Training
The idea that one should be able to prove that a particular training process took place—rather than merely report a hash of the final model—was articulated by the Proof-of-Learning (PoL) agenda, which formalized the notion of training trajectory evidence and its threat models [1]. Subsequent analyses refined the economic and strategic dimensions, arguing that proofs must remain sound even when parties have incentives to game audits, thus introducing incentive-compatible variants of PoL and clarifying adversarial capabilities around partial logging and cherry-picked checkpoints [19]. At the same time, the limits literature has tempered expectations about what can be established a posteriori: information-theoretic and statistical arguments show that membership inference can neither reliably confirm nor refute that a specific record was used in training, particularly for large overparameterized models [23, 24]. These observations do not negate PoL; rather, they motivate pre-commitment mechanisms where the trainer binds themselves to the trajectory and data before the fact, a stance we inherit in VFT.
@@
Verifiable Machine Learning (VML) takes a more computational route by transporting training or inference into succinct proof systems. Early milestones demonstrated that nontrivial neural computations can be encoded with acceptable overheads, first for inference and accuracy claims over vision models [3] and then for broader neural operators via efficient arithmetization and conversion frameworks [5]. A line of systems work explored zero-knowledge inference pipelines with practical prover and verifier runtimes, showing that privacy-preserving verification of predictions can be integrated into ML services [15]. Lifting these guarantees to training is more challenging—gradients, optimizer states, and stochastic sampling substantially enlarge circuits—but recent systems make tangible progress: zkDL proposes efficient proofs for deep learning training by exploiting structure and approximations [9], Kaizen develops zero-knowledge proofs specifically for training deep networks with attention to optimizer semantics [11], and zkLLM targets language models by leveraging proof-friendly tensor factorizations and approximation to softmax and normalization layers [13]. Complementary advances in proof theory, such as series-expansion techniques for non-linearities and careful error accounting in the arithmetization of common layers, further reduce prover costs and extend coverage to richer model families [14].
@@
Beyond single systems, surveys and position papers map the design space and sharpen research questions. Overviews of zero-knowledge ML outline core primitives (commitments, lookup arguments, sum-check protocols), recurring engineering bottlenecks (non-linearities, memory), and deployment pathways (off-chain aggregation, on-chain verification) [6, 8]. A focused look at the on-chain dimension connects proof-friendly ML to distributed ledgers, highlighting programmability and cost models when verifiable computation must feed on-chain governance or marketplaces [12]. Our work adopts a hybrid posture: proofs are generated off-line and aggregated, while logs and key commitments remain auditable and, when appropriate, registrable.
@@ II-B Data and Pipeline Provenance
Verification of training is hollow without credible claims about the data. The Data Provenance Initiative documented licensing and attribution issues across widely used corpora, demonstrating how ambiguous or missing metadata undermine compliance, reproducibility, and ethics [2]. At a complementary level, tools for verifying training data investigate whether one can efficiently check that certain records (or distributions of records) must have been present, even when raw datasets cannot be shared [22]. Distributional attestation pushes in a similar direction, proposing cryptographic protocols to certify statistical properties of training data without revealing examples, so that downstream verifiers at least know the shape of what was ingested [7].
A parallel systems thread treats provenance as a first-class artifact in ML engineering. Automated screening over data preparation pipelines can flag risks and policy violations early, leveraging provenance graphs to surface transformations and their implications [4]. Practical capture of end-to-end provenance across experimentation frameworks provides the raw substrate for audit trails and reproducible science; MLflow2PROV, for example, extracts structured lineage from common ML tooling with minimal burden on practitioners [10]. Building on such capture, recent work argues for integrating provenance into the operational life-cycle—what is recorded, how it is queried, and how it is bound to released artifacts—so that downstream consumers can validate claims about data, configurations, and code versions [35]. When metadata are incomplete, even model-assisted inference of provenance may help recover partial lineage, though such approaches must be treated with caution given the risk of hallucination or confirmation bias [25]. We view these tools as complementary to VFT: provenance can anchor which datasets were eligible, and proofs can attest which batches and updates were actually used.
@@ II-C Output Attribution via Watermarking and Its Limits
Watermarks are attractive because they are lightweight and retrofittable. Early methods for LLM text embed pseudo-random beacons that can be detected statistically without changing human-perceived fluency [26]. Large-scale evaluations suggest that scalable detectors can operate at Internet scale, though operating points depend sensitively on sampling and domain shifts [27]. For images, invisible but robust fingerprinting for diffusion models extends the concept to the visual domain [33]. Yet the reliability of such signals is contested: formal analyses show regimes where watermarks are learnable by downstream models or where adversaries can manipulate generation to evade detection [28], while empirical attacks demonstrate removal and even theft of watermark keys [29, 32]. This tension has spurred coding-theoretic and cryptographic redesigns that seek provable robustness guarantees under realistic edits [30, 31], as well as surveys that summarize the trade space and caveats for deployment [34]. In our scope, watermarking serves as an output-side signal that complements, but cannot replace, training-time verifiability.
@@ II-D Hardware-Assisted Attestation and ML Property Cards
Trusted execution environments (TEEs) and attestation mechanisms promise stronger integrity for training and deployment pipelines by anchoring code identity and configuration in hardware-rooted proofs. Recent work moves beyond attest-or-prove dichotomies and advocates verifiable property cards: structured attestations about what was trained and how, designed to be composable with cryptographic verification and external audits [21]. In practice, TEEs can justify sensitive steps (e.g., decrypting private data, enforcing configuration constraints), but evidence of correct gradient updates and data usage still benefits from VML-style proof artifacts. We therefore treat TEEs as trustworthy enforcers around data access and hyperparameter policy, and ZK proofs as trustworthy witnesses of the numerical computations themselves.

### v2

@@ I Introduction
Large language models (LLMs) have rapidly migrated from research prototypes to critical infrastructure for search, assistance, and decision support, yet the trust gap around how these models are trained and fine-tuned is widening. Concerns span from whether proprietary or sensitive data were used, to whether claimed training configurations actually produced the released checkpoints, and to whether post-hoc forensics can establish accountability when things go wrong [longpre2024data]. Conventional privacy and security analyses underscore that inferring whether specific records participated in training is fundamentally limited, which makes ex post verification particularly tenuous [hu2022miaSurvey, zhang2024miaProofs]. Proof-of-learning (PoL) pushes the opposite direction by seeking cryptographic or audit-style evidence that the training trajectory is genuine [jia2021proof], and recent variants address incentive compatibility when proofs drive economic or governance consequences [zhao2024polIncentive].
@@
A parallel thread pursues verifiable machine learning (VML), where the computation itself is accompanied by succinct proofs. Early systems established feasibility for neural inference under zero-knowledge (ZK) or related proof systems [liu2021zkcnn, weng2021mystique, wang2023ezdps]. More recently, researchers have begun to lift verification from inference to (parts of) training by circuitizing differentiable updates and optimizer states [sun2023zkdl, abbaszadeh2024zkpot]. For LLMs, architectural regularities and approximation-friendly algebra enable practical proof composition, suggesting a path toward end-to-end guarantees in realistic instruction-tuning pipelines [sun2024zkllm]. Methodologically, advances in proof-friendly tensor approximations (e.g., polynomial constraints over softmax) and series-expansion techniques further reduce prover overheads and widen coverage [taylorZKML2025]. Survey and position papers consolidate these advances and highlight open challenges around scalability, recursion, and on-chain verification for ML [kersic2024onchainzkml, zkvml2023survey, peng2025zkmlsurvey].
@@
Beyond cryptographic protocols, the surrounding audit stack must establish that the data and configurations claimed for training are accurate. Dataset-level attestation proposes to certify distributional properties without revealing examples [dold2024attesting], while data-engineering provenance offers pragmatic hooks by capturing end-to-end lineage across preparation and experimentation workflows [schelter2024provenanceScreen, schlegel2023mlflow2prov]. Systems work is beginning to close the loop by treating provenance as a first-class runtime artifact amenable to programmatic enforcement and reproducibility [schlegel2025isProvenance], and by exploring how LLMs themselves can assist in provenance inference when metadata are incomplete [almuntashiri2025inferProv]. Complementarily, model-provenance testing proposes behavioral probes to ascribe models to sources or families when direct logs are unavailable [nikolic2025provenance]. Together with PoL-style arguments [jia2021proof, zhao2024polIncentive], these strands motivate holistic, multi-layer verification.
@@
Watermarking techniques have also been advanced as a lightweight provenance signal for model outputs, but their reliability is contested. Classic and modern schemes embed statistical beacons into generated text and images [kirchenbauer2023watermark, wen2023treering], with recent work demonstrating scalable detectors for LLM text at industrial scale [dathathri2024natureWatermark]. Yet watermarks can be fragile or learnable by downstream models [gu2024learnability], susceptible to removal or theft [jovanovic2024wmstealing, zhang2024mipSteal], and often deliver weakest guarantees precisely where high-stakes provenance is needed, prompting more robust coding-theoretic designs [zhao2023provable, qu2025multibit] and broader surveys that map the design space and limitations [liu2024textwmSurvey]. These observations reinforce the case for training-time verifiability that does not rely solely on output beacons.
@@
In parallel, trusted execution and attestation are being used to produce verifiable “property cards” about ML artifacts, offering complementary hardware-backed assurances when cryptographic proofs are too costly or incomplete [duddu2025laminator]. However, TEEs alone cannot resolve disputes about which data were used or whether the optimizer trajectory conformed to policy, again pointing toward hybrid designs that combine attestations, provenance capture, and cryptographic proofs.
@@
The need for verifiable training is especially acute in federated and edge-centric learning, where heterogeneous clients, intermittent connectivity, and privacy constraints complicate governance. Practical systems already struggle with non-IID data, communication bottlenecks, and asynchrony [liu2021communication, liu2021adaptive, wang2021resource]. Algorithmic advances target adaptive regularization and knowledge distillation for robustness [liu2023adaptive], decentralized aggregation with layer-wise strategies [liu2023yoga], and personalization frameworks that better reflect client idiosyncrasies [yao2024ferrari, yan2024peaches]. Neural architecture search and progressive training mitigate heterogeneity in on-device settings [10251628, liu2024enhancing], while LoRA- and quantization-aware fine-tuning reduce resource footprints for practical deployments [li2025fedquad]. Recent work explores probabilistic communication and local-update strategies to accelerate decentralized training under diverse edges [liu2025accelerating, liu2025adaptive], pipelines for collaborative end–cloud inference [gao2025accelerating], strategies for MoE inference efficiency [yan2025accelerating], communication-efficient decentralized graph learning [wang2025towards], and mitigating catastrophic forgetting during federated fine-tuning [huo2025mitigating]. These directions improve scalability but simultaneously enlarge the attack surface and the number of untrusted actors—prime territory for verifiable fine-tuning protocols that certify data usage and update correctness [schlegel2025isProvenance, zkvml2023survey].
@@
Ubiquitous sensing further complicates the risk calculus. Wi-Fi and RF modalities, often intertwined with intelligent assistants, have been shown to support keystroke eavesdropping from commodity microphones [10587029], undetectable attacks on PHY-layer fingerprint-based authentication [10339891], and open-set gesture and attention-based recognition that blur lines between ambient intelligence and surveillance [10899398, 10090421]. Multimodal affective computing demonstrates how Wi-Fi CSI and vision together can unobtrusively infer emotion, expanding the privacy stakes of data pipelines [10149418]. On the defensive and robust-sensing side, prior work models and mitigates co-channel interference in activity recognition through subcarrier correlation selection and interference-independent phase components [9075376, 9613773]. Healthcare-oriented Wi-Fi sensing shows both potential and sensitivity, from target-oriented respiratory monitoring to pulmonary function estimation without physical contact [10804189, 10699348]. RF-based writing and identity inference illustrates how ordinary RFID can leak highly personal behaviors [feng2025rf]. Even adjacent work on label-noise suppression and distributionally robust facial expression recognition underscores how data curation and reliability shape downstream behaviors that later demand auditing [10916982, cui2025learningheterogeneitygeneralizingdynamic]. These examples motivate verifiable training pipelines that can convincingly demonstrate which signals were included, which were excluded, and how updates adhered to declared policies [schelter2024provenanceScreen].
@@
Our perspective in this paper is that verifiable fine-tuning should be treated as a first-class systems problem that composes three pillars: (i) cryptographic protocols that prove the existence of a legitimate gradient-based path from a public initialization to the released model under declared hyperparameters [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], (ii) provenance capture that binds those proofs to accountable, inspectable data and configuration claims [schlegel2023mlflow2prov, schlegel2025isProvenance, nikolic2025provenance], and (iii) audit mechanisms that recognize the limits of post-hoc inference and instead emphasize pre-commitment and incentive-compatible verification [jia2021proof, fang2022limits, zhao2024polIncentive]. We position our approach within this emerging landscape and focus on end-to-end verifiability for fine-tuning, complementing watermarking-based output attribution [kirchenbauer2023watermark, dathathri2024natureWatermark] and hardware-assisted attestations [duddu2025laminator], while remaining deployable in federated and edge scenarios [10251628, yao2024ferrari].
@@ II-A From Proof-of-Learning to Verifiable Training
The idea that one should be able to prove that a particular training process took place—rather than merely report a hash of the final model—was articulated by the Proof-of-Learning (PoL) agenda, which formalized the notion of training trajectory evidence and its threat models [jia2021proof]. Subsequent analyses refined the economic and strategic dimensions, arguing that proofs must remain sound even when parties have incentives to game audits, thus introducing incentive-compatible variants of PoL and clarifying adversarial capabilities around partial logging and cherry-picked checkpoints [zhao2024polIncentive]. At the same time, the limits literature has tempered expectations about what can be established a posteriori: information-theoretic and statistical arguments show that membership inference can neither reliably confirm nor refute that a specific record was used in training, particularly for large overparameterized models [hu2022miaSurvey, zhang2024miaProofs]. These observations do not negate PoL; rather, they motivate pre-commitment mechanisms where the trainer binds themselves to the trajectory and data before the fact, a stance we inherit in VFT.
@@
Verifiable Machine Learning (VML) takes a more computational route by transporting training or inference into succinct proof systems. Early milestones demonstrated that nontrivial neural computations can be encoded with acceptable overheads, first for inference and accuracy claims over vision models [liu2021zkcnn] and then for broader neural operators via efficient arithmetization and conversion frameworks [weng2021mystique]. A line of systems work explored zero-knowledge inference pipelines with practical prover and verifier runtimes, showing that privacy-preserving verification of predictions can be integrated into ML services [wang2023ezdps]. Lifting these guarantees to training is more challenging—gradients, optimizer states, and stochastic sampling substantially enlarge circuits—but recent systems make tangible progress: zkDL proposes efficient proofs for deep learning training by exploiting structure and approximations [sun2023zkdl], Kaizen develops zero-knowledge proofs specifically for training deep networks with attention to optimizer semantics [abbaszadeh2024zkpot], and zkLLM targets language models by leveraging proof-friendly tensor factorizations and approximation to softmax and normalization layers [sun2024zkllm]. Complementary advances in proof theory, such as series-expansion techniques for non-linearities and careful error accounting in the arithmetization of common layers, further reduce prover costs and extend coverage to richer model families [taylorZKML2025].
@@
Beyond single systems, surveys and position papers map the design space and sharpen research questions. Overviews of zero-knowledge ML outline core primitives (commitments, lookup arguments, sum-check protocols), recurring engineering bottlenecks (non-linearities, memory), and deployment pathways (off-chain aggregation, on-chain verification) [zkvml2023survey, peng2025zkmlsurvey]. A focused look at the on-chain dimension connects proof-friendly ML to distributed ledgers, highlighting programmability and cost models when verifiable computation must feed on-chain governance or marketplaces [kersic2024onchainzkml]. Our work adopts a hybrid posture: proofs are generated off-line and aggregated, while logs and key commitments remain auditable and, when appropriate, registrable.
@@ II-B Data and Pipeline Provenance
Verification of training is hollow without credible claims about the data. The Data Provenance Initiative documented licensing and attribution issues across widely used corpora, demonstrating how ambiguous or missing metadata undermine compliance, reproducibility, and ethics [longpre2024data]. At a complementary level, tools for verifying training data investigate whether one can efficiently check that certain records (or distributions of records) must have been present, even when raw datasets cannot be shared [choi2023tools]. Distributional attestation pushes in a similar direction, proposing cryptographic protocols to certify statistical properties of training data without revealing examples, so that downstream verifiers at least know the shape of what was ingested [dold2024attesting].
A parallel systems thread treats provenance as a first-class artifact in ML engineering. Automated screening over data preparation pipelines can flag risks and policy violations early, leveraging provenance graphs to surface transformations and their implications [schelter2024provenanceScreen]. Practical capture of end-to-end provenance across experimentation frameworks provides the raw substrate for audit trails and reproducible science; MLflow2PROV, for example, extracts structured lineage from common ML tooling with minimal burden on practitioners [schlegel2023mlflow2prov]. Building on such capture, recent work argues for integrating provenance into the operational life-cycle—what is recorded, how it is queried, and how it is bound to released artifacts—so that downstream consumers can validate claims about data, configurations, and code versions [schlegel2025isProvenance]. When metadata are incomplete, even model-assisted inference of provenance may help recover partial lineage, though such approaches must be treated with caution given the risk of hallucination or confirmation bias [almuntashiri2025inferProv]. We view these tools as complementary to VFT: provenance can anchor which datasets were eligible, and proofs can attest which batches and updates were actually used.
@@ II-C Output Attribution via Watermarking and Its Limits
Watermarks are attractive because they are lightweight and retrofittable. Early methods for LLM text embed pseudo-random beacons that can be detected statistically without changing human-perceived fluency [kirchenbauer2023watermark]. Large-scale evaluations suggest that scalable detectors can operate at Internet scale, though operating points depend sensitively on sampling and domain shifts [dathathri2024natureWatermark]. For images, invisible but robust fingerprinting for diffusion models extends the concept to the visual domain [wen2023treering]. Yet the reliability of such signals is contested: formal analyses show regimes where watermarks are learnable by downstream models or where adversaries can manipulate generation to evade detection [gu2024learnability], while empirical attacks demonstrate removal and even theft of watermark keys [jovanovic2024wmstealing, zhang2024mipSteal]. This tension has spurred coding-theoretic and cryptographic redesigns that seek provable robustness guarantees under realistic edits [zhao2023provable, qu2025multibit], as well as surveys that summarize the trade space and caveats for deployment [liu2024textwmSurvey]. In our scope, watermarking serves as an output-side signal that complements, but cannot replace, training-time verifiability.
@@ II-D Hardware-Assisted Attestation and ML Property Cards
Trusted execution environments (TEEs) and attestation mechanisms promise stronger integrity for training and deployment pipelines by anchoring code identity and configuration in hardware-rooted proofs. Recent work moves beyond attest-or-prove dichotomies and advocates verifiable property cards: structured attestations about what was trained and how, designed to be composable with cryptographic verification and external audits [duddu2025laminator]. In practice, TEEs can justify sensitive steps (e.g., decrypting private data, enforcing configuration constraints), but evidence of correct gradient updates and data usage still benefits from VML-style proof artifacts. We therefore treat TEEs as trustworthy enforcers around data access and hyperparameter policy, and ZK proofs as trustworthy witnesses of the numerical computations themselves.